Two capacitors C1 and C2 of 1μf and 3μf are connected in parallel

across a voltage supply. At a particular instant when the applied

voltage is increasing positively, the two capacitors together draw a

current of 4A. Find the instantaneous rate of increase of potential

across them. 


1
Expert's answer
2021-10-04T04:55:30-0400

Given

C1= 1uF

C2= 3uF

are connected in parallel

Total capacitance C= C1+C2=(1+3)uF=4 uF

"dV\/dt= I\/C"

= 4* 10-6/4=1*10-6

"dV\/dt=1"0-6 volt/time